OpenStudy (anonymous):

a square carpet covers 75% of the area of a floor. the floor is 8 m by 8 m. What are the dimensions of the carpet? give ur answer to two decimal places. what area of the floor is not covered by the carpet?

OpenStudy (anonymous):

area of the floor = 8 x 8 = 64m^2 3/4 (64) = 48 m^2 dimensions of the carpet = \[\sqrt{48}\] = \[4(\sqrt{3})\]

OpenStudy (anonymous):

thanks but what does ^ mean again?